# Memorization to Generalization: Emergence of Diffusion Models from Associative Memory

### v2

@@ E.3 MNIST
Appendix F Additional Visualizations of the Three Sample Types
F.1 Memorized Samples
@@
Figure 15: Visualizations of additional memorized patterns and their top-4 nearest neighbors for different datasets. The top row illustrates nearest neighbors from the training set while the bottom row depicts those from the synthetic set. Memorized samples are duplicates of the training set. During the strong memorization phase, duplicates are also found within the synthetic set. Note, even though our memorized detection metric (Eq. 13) does not utilize the synthetic set, we are showing the nearest neighbors obtained from it, for consistency.
F.2 Spurious Samples
@@
Figure 16: Visualizations of additional spurious patterns and their top-4 nearest neighbors for different datasets. The top row depicts nearest neighbors from the training set, while the bottom row shows those from the synthetic set. Spurious patterns are demonstrated to arise from the onset of generalization where the mixing of training data points begins. Since the model’s generalization is at its infancy, duplicates of spurious patterns sometimes appear several times in the synthetic set $S^{\prime}$, much like memorized patterns. Additionally, these samples lack the uniqueness to be considered as generalized samples as the model has yet to fully learn the underlying data distribution.
F.3 Generalized Samples
@@
Figure 17: Visualization of generalized patterns and their top-4 nearest neighbors for different datasets. The top row illustrates nearest neighbors from the training set while the bottom row depicts those from the synthetic set. Generalized samples are novel samples, which have little to no resemblance to their nearest neighbors in training and synthetic sets.
Appendix G Experimental Details on the Basin of Attraction Analysis Across the Transition
@@
Algorithm Details. To obtain the results recorded in the above Figs. (18) and (19), we first computed the critical time $t_{c}$ given an image ${\mathbf{x}}_{0}={\mathbf{y}}$. Specifically, this is done by using Algorithm (1), where $t_{c}$ is obtained using the deterministic DDIM [24] sampler to denoise the perturbed version ${\mathbf{x}}_{t}$ of ${\mathbf{x}}_{0}$ for each step in $t=1,\dots,T$, given $N^{\prime}$ number of denoising steps starting backward from $t$ to 00. To save computational time, the algorithm utilizes a stride constant $s$ to skip over some of the timesteps $t=1,1+s,1+2s,\dots,T-s-1$. For each timestep $t$, the variance preserving forward process is applied onto the targeted image ${\mathbf{x}}_{0}$ yielding ${\mathbf{x}}_{t}$. To account for the variations in noise, the image ${\mathbf{x}}_{0}$ is perturbed for $M$ number of times at each step $t$.
To check whether if the critical time $t_{c}$ is reached, the algorithm utilizes a distance metric $d$ — which is LPIPS [77] equipped with the vgg-16 backbone [82] for all of the datasets — checking if a recovered image $\hat{{\mathbf{x}}}_{0}$ obtained from the DDIM sampler [24] with $N^{\prime}$ denoising steps is similar to the original pattern ${\mathbf{x}}_{0}$ within an error threshold $\delta_{d}\in(0,1)$. Unlike the experiment on the memorization-generalization transition detailed in Section 5, we can afford to trade computational time for better accuracy in this experiment by using a larger backbone for LPIPS. Please note that the metric LPIPS is bounded between 0 (most similar) and 1 (least similar). To keep track of the critical time $t_{c}$ for all perturbations, we allow it to be a vector of integers $t_{c}\in\mathbb{N}^{M}$, where $M$ is the number of perturbations or trials for the image ${\mathbf{x}}_{0}$. To ensure that $t_{c}$ is correct, we have a binary stopping vector $\hat{S}\in\{0,1\}^{M}$, which keeps track of whether each perturbation of ${\mathbf{x}}_{0}$ is no longer recoverable, such that we only update an entry of the vector $t_{c}$ if and only if the corresponding entry in $\hat{S}$ is a zero (not stopped).
@@
Since the entire process of finding $t_{c}$ is a rough search over a set of timesteps, we require a “breaking condition” for the search. In Algorithm 1, this condition is determined by the value $p\in[0,1]$, which is a probability corresponding to the recovery probability of the image ${\mathbf{x}}_{0}$ with respect to its $M$ perturbations. To compute $p$, we calculate the distance vector $\mathbf{d}=d(\hat{{\mathbf{x}}}_{0},{\mathbf{x}}_{0})$ where $\hat{{\mathbf{x}}}_{0}$ is the predicted version of ${\mathbf{x}}_{0}$ obtained from denoising ${\mathbf{x}}_{t}$. Then, we compute the recovery binary vector $\mathbf{r}\in[0,1]^{M}$ by applying the indicator function onto $\mathbf{d}$ using the threshold $\delta_{d}$. Lastly, we calculate the probability value $p=\frac{1}{M}\sum^{M}_{i=1}\mathbf{r}_{i}$. If this value $p$ is below the probability threshold $\delta_{p}$, we stop the search for critical time across all trials.
Experimentation Details. For each training data size $K$, we used a maximum of the top 512 samples belonging to each of the sample types — e.g., training data, memorized, spurious, and generalized samples — to find their critical time $t_{c}$ using Algorithm 1. Note, with the exception of the training data points, we sorted samples belonging to each set — identified via the detection metrics (13), (14) and (15) — by their distance from least to greatest. We computed the log volume using the following equation
$V({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}}=\frac{\pi^{\frac{N}{2}}}{\Gamma(\frac{N}{2}+1)}R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})^{N}$ (33)
where $\Gamma(\cdot)$ is the Euler’s gamma function [83] and $N$ is the dimensionality of the image. The radius $R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})$ is the Euclidean norm of the difference between the original image ${\mathbf{x}}_{0}$ and its recoverable perturbation ${\mathbf{x}}_{t_{c}}$. For Figs. (6), (18), and (19), we plotted the average log-volume value per each training data size $K$. Meanwhile, for Algorithm 1, we fixed the stopping probability $\delta_{p}$ as $0.8$, the number of perturbations $M$ as $20$, the DDIM steps $N^{\prime}$ as $10$ and the stride constant $s$ as $10$ for all datasets. At some training data size $K$, in which there are very few samples in either memorized, spurious or generalized set, we opted to not include the statistics of such a set, partly due to its large variance. We remove the statistics of such a set if the number of its samples is less than $0.1\%$ of its corresponding synthetic set for a training data size $K$.
For CIFAR10 and LSUN-CHURCH, we utilized the threshold $\delta_{d}$ as $0.03$ and $0.12$ respectively. We kept these threshold values similar to the memorized threshold values $\delta_{m}$ detailed in Table 2, since they have proven to be effective at separating the memorized samples from non-memorized ones in the previous experiment. However, since it is difficult to select the threshold value for the $L_{2}$-distance metric — because the range of the metric is between 00 and $\infty$ — we opted to use LPIPS to experiment with MNIST and FASHION-MNIST using the same threshold $\delta_{d}=0.1$ for both. To find this optimal threshold value, we selected three data sizes, where each belongs to one of the three phases: memorization, spurious, or generalization. We then observed which of the several threshold values, starting from $0.01$ to $0.15$ with the increment of $0.01$, reflects best of the performance given by LPIPS using vgg-16 [82] as its backbone. Specifically, we evaluated these thresholds using $K\in\{4000,11724,60000\}$ for MNIST, and $K\in\{4000,8965,60000\}$ for FASHION-MNIST. We simply perform Algorithm 1 for those three data sizes on a small batch of $32$ samples for the three sample types, and manually check whether the recovered samples (using the obtained critical time $t_{c}$ to denoise from) are similar to the original images and if this observation is consistent across those three data sizes.
G.1 Algorithm for Computing the Critical Time
1
2
3Inputs
4        Pattern ${\mathbf{y}}$
5        Score model $s_{\theta}$
6        DDIM sampler $\Phi$
7        Error threshold $\delta_{d}$
8        Distance metric $d$
9        Denoising steps $N^{\prime}$
10        Total diffusion time $T$
11        Time stride $s$
12        Number of trials $M$
13        Probability threshold $\delta_{p}$
14        Cumulative variances $\bar{\alpha}$
15        Variances $\beta$
16Compute $t_{c}$
        ${\mathbf{x}}_{0}\leftarrow\text{duplicate}({\mathbf{y}},M)$
         // Duplicate pattern ${\mathbf{y}}$ for $M$ times
        $t_{c}\leftarrow\mathbf{0}$
@@ 15        Variances $\beta$
17        for $t=1,1+s,1+2s,\dots,T-s-1$ do
               ${\mathbf{z}}_{t}\sim\mathcal{N}(0,\mathbf{I})$
@@
18
              ${\mathbf{x}}_{t}\leftarrow\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\mathbf{z}}_{t}$
                // Forward diffusion
19
              $\hat{{\mathbf{x}}}_{0}\leftarrow\Phi(s_{\theta},{\mathbf{x}}_{t},t,\bar{\alpha},\beta,N^{\prime})$
@@
23
24              if $p<\delta_{p}$ then
25                      return $t_{c}$
26              else
27                      $\hat{S}\leftarrow\text{update-stop}(\hat{S},\mathbf{r})$
28                      $t_{c}\leftarrow\text{update-time}(t,t_{c},\hat{S})$
29
30
31Output
32        Critical time $t_{c}\in\mathbb{N}^{M}$
Algorithm 1 Compute Critical Time $t_{c}$ for a single pattern ${\mathbf{y}}$
G.2 The Basin of Attraction Across the Transition for All Datasets
@@
Figure 18: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in CIFAR10 and LSUN-CHURCH computed using Equation 33 with up to 512 samples per type. When fewer than 512 samples are available, all points are used. The log-volume of training samples and the transition plot from Figure 5 are included to confirm the shrinking basins of attraction in diffusion models. Variations in the critical time $t_{c}$ during the transition, which affect the radius $R$, lead to distinct log-volume trends across sample types, see the recorded critical times in Figure 20. Shaded regions indicate standard deviation of the log-volume.
@@
Figure 19: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in FASHION-MNIST and MNIST, computed using Equation 33 with up to 512 samples per type. When fewer than 512 samples are available, all points are used. The log-volume of training samples and the transition plot from Figure 5 are included to confirm the shrinking basins of attraction in diffusion models. Variations in critical time $t_{c}$ and radius $R$ lead to distinct log-volume trends across sample types, see the recorded critical times in Figure 21. Shaded regions indicate standard deviation of the log-volume.
G.3 Average Critical Time Across The Transition
@@
Figure 20: Average critical time, obtained using Algorithm 1, with a maximum of 512 samples across the memorization-generalization transition for each sample type in CIFAR10 and LSUN-CHURCH. The average critical times for the training data samples at each training data size $K$ are included to illustrate the overall decrease in the time and also the amount of perturbation in which the diffusion model can effectively recover the original sample within some error. Each sample type exhibits a distinctive critical time at each training data size $K$. The shaded region (or the error bars) shown here is the standard deviation computed from all of the critical times at each $K$ with respect to each type of samples. The transition plot of each dataset is shown here to visually guide where the memorization-generalization transition occurs.
@@
Figure 21: Average critical time, obtained using Algorithm 1, with a maximum of 512 samples across the memorization-generalization transition for each sample type in FASHION-MNIST and MNIST. The average critical times for the training data samples at each training data size $K$ are included to illustrate the overall decrease in the time and also the amount of perturbation in which the diffusion model can effectively recover the original sample within some error. Each sample type exhibits a distinctive critical time at each training data size $K$. The shaded region (or the error bars) shown here is the standard deviation computed from all of the critical times at each $K$ with respect to each type of samples. The transition plot of each dataset is shown here to visually guide where the memorization-generalization transition occurs.
G.4 Visual Demonstration of the Shrinkage of the Basins of Attraction
Experiment Details. For Figure 7, we randomly selected a subset of training images that are shared across all 38 diffusion models, each trained with a different training set size. To determine an appropriate timestep $t\approx t_{c}$, we referred to Figure 20, which plots the average critical time $t_{c}$ as a function of training data size $K$. From this plot, we identified a plateau point in the memorization fraction curve, occurring around $K=3931$, and selected $t_{c}\approx 450$ for CIFAR10. Using the timestep starting at $450$, each image was perturbed via the forward process and then denoised using the DDIM sampler [24] with 25 steps across all models (starting backward from the indicated timestep). We manually examined the denoised outputs across training sizes to assess whether the resulting generations effectively illustrate the shrinkage of the basin of attraction. If the output appeared overly memorized (or too close to the original), we increased the timestep by increments of 10 and repeated the process to obtain a more illustrative transition. Lastly, in the case of LSUN-CHURCH, we selected the initial timestep to be $550$ accordingly to Figure 20.
Appendix H Additional Details on the Relative Energy Computation and its Experimentation
In this section, we provide details on the formulation of the relative energy done by [38] and an overview on the experiment done in Section 7 of the main text.
Relative Energy Formulation. Consider the SDE for the reverse process in Equation 2 of the main text
$\mathrm{d}{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})]\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (34)
We can re-express the above equation with the potential function $u(x,t)$
$-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)=\mathbf{f}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ (35)

### v3

@@ B.6  Additional Visualizations of Distance Histogram
B.7  Additional Examples of the Three Distinct Sample Types
@@
Figure 13: Visualizations of additional memorized patterns and their top-4 nearest neighbors for different datasets. The top row illustrates nearest neighbors from the training set while the bottom row depicts those from the synthetic set. Memorized samples are duplicates of the training set $\mathsf{S}$. During the strong memorization phase, duplicates are also found within the synthetic set $\mathsf{S}^{\prime}$. Note, even though our memorized detection metric does not utilize the synthetic set, we are showing the nearest neighbors obtained from it, for consistency.
@@
Figure 14: Visualizations of additional spurious patterns and their top-4 nearest neighbors for different datasets. The top row depicts nearest neighbors from the training set, while the bottom row shows those from the synthetic set. Spurious patterns are demonstrated to arise from the onset of generalization where the mixing of training data points begins. Since the model’s generalization is at its infancy, duplicates of spurious patterns sometimes appear several times in the synthetic set $\mathsf{S}^{\prime}$, much like memorized patterns. Additionally, these samples lack the uniqueness to be considered as generalized samples as the model has yet to fully learn the underlying data distribution.
@@
Figure 15: Visualization of generalized patterns and their top-4 nearest neighbors for different datasets. The top row illustrates nearest neighbors from the training set $\mathsf{S}$ while the bottom row depicts those from the synthetic set $\mathsf{S}^{\prime}$. Generalized samples are novel samples, which have little to no resemblance to their nearest neighbors in training and synthetic sets.
Appendix C The Volume of the Basin of Attraction
C.1  Algorithm Details
@@
To obtain the average log-volume results recorded in fig.˜4 of the main text, we first computed the critical time $t_{c}$ given an image $\hat{{\mathbf{x}}}={\mathbf{x}}_{0}$. Specifically, this is done by using Alg. (1), where $t_{c}$ is obtained using the deterministic DDIM [60] sampler to denoise the perturbed version ${\mathbf{x}}_{t}$ of ${\mathbf{x}}_{0}$ for each step in $t=1,\dots,T$, given $N^{\prime}$ number of denoising steps starting backward from $t$ to 0. To save computational time, the algorithm utilizes a stride constant $s$ to skip over some of the timesteps $t=1,1+s,1+2s,\dots,T-s-1$. For each timestep $t$, the variance preserving forward process is applied onto the targeted image ${\mathbf{x}}_{0}$ yielding ${\mathbf{x}}_{t}$. To account for the variations in noise, the image ${\mathbf{x}}_{0}$ is perturbed for $M$ number of times at each step $t$.
[…]
@@
Since the entire process of finding $t_{c}$ is a rough search over a set of timesteps, we require a ‘breaking condition’ for the search. In Alg. (1), this condition is determined by the value $p\in[0,1]$, which is a probability corresponding to the recovery probability of the image ${\mathbf{x}}_{0}$ with respect to its $M$ perturbations. To compute $p$, we calculated the distance vector $\mathbf{d}=d(\hat{{\mathbf{x}}}_{0},{\mathbf{x}}_{0})$ where $\hat{{\mathbf{x}}}_{0}$ is the predicted version of ${\mathbf{x}}_{0}$ obtained from denoising ${\mathbf{x}}_{t}$. Then, we computed the recovery binary vector $\mathbf{r}\in[0,1]^{M}$ by applying the indicator function onto $\mathbf{d}$ using the threshold $\delta_{d}$. Lastly, we calculated the probability value $p=\frac{1}{M}\sum^{M}_{i=1}\mathbf{r}_{i}$. If this value $p$ is below the probability threshold $\delta_{p}$, we stop the search for critical time across all trials.
Inputs
  1.5pt Pattern $\hat{{\mathbf{x}}}$
  1.5pt Time stride $s$
  1.5pt Score model $s_{\theta}$
  1.5pt DDIM sampler $\Phi$
  1.5pt Error threshold $\delta_{d}$
  1.5pt Distance metric $d$
  1.5pt Denoising steps $N^{\prime}$
  1.5pt Total diffusion time $T$
  1.5pt Number of trials $M$
  1.5pt Probability threshold $\delta_{p}$
  1.5pt Forward process $\mathcal{F}$
end
Compute $t_{c}$
  5pt ${\mathbf{x}}_{0}\leftarrow\text{duplicate}(\hat{{\mathbf{x}}},M)$
 // Duplicate pattern for $M$ times
  5pt $t_{c}\leftarrow\mathbf{0}$
@@
  5pt for $t=1,1+s,1+2s,\dots,T-s-1$ do
     5pt $\epsilon\sim\mathcal{N}(0,I)$
@@
     5pt
    ${\mathbf{x}}_{t}\leftarrow\mathcal{F}({\mathbf{x}}_{0},\epsilon,t)$
    // Forward process
     5pt
    $\hat{{\mathbf{x}}}_{0}\leftarrow\Phi(s_{\theta},{\mathbf{x}}_{t},t,N^{\prime})$
@@
     8pt
    if $p<\delta_{p}$ then
       return $t_{c}$
    else
       $\hat{S}\leftarrow\text{update}(\hat{S},\mathbf{r})$
[…]
      end if
   end for
end
Output
  5pt Critical time $t_{c}\in\mathbb{N}^{M}$
end
Algorithm 1 Compute critical time $t_{c}$ for a pattern
C.2  Experimental Details
For each training data size $K$, we used a maximum of the top 512 samples belonging to each of the sample types: training data, memorized, spurious, and generalized samples. Note, with the exception of the training data points, we sorted samples belonging to each set, identified via our detection metrics, by their distance from least to greatest. We computed the logarithmic version of the following $N$-ball volume equation:
$V(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})=\frac{\pi^{\frac{N}{2}}}{\Gamma(\frac{N}{2}+1)}R(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})^{N}$ (34)
using the $t_{c}$ we computed from Alg. (1) where $\Gamma(\cdot)$ is the Euler’s gamma function [55] and $N$ is the dimensionality of the image. The radius $R(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})$ is the Euclidean norm of the difference between a given image $\hat{{\mathbf{x}}}\in\mathbb{R}^{N}$ and its recoverable perturbation ${\mathbf{x}}_{t_{c}}$ computed from the forward process. For Figs. (16)-(17), we plotted the average log-volume value calculated for each training data size $K$ per dataset.
Meanwhile, for Alg. (1), we fixed the stopping probability $\delta_{p}$ as $0.8$, the number of perturbations $M$ as $20$, the DDIM steps $N^{\prime}$ as $10$ and the stride constant $s$ as $10$ for all datasets. At some training data size $K$, in which there are very few samples in either memorized, spurious or generalized set, we opted to not include the statistics of such a set, partly due to its large variance. We remove the statistics of such a set if the number of its samples is less than $0.1\%$ of its corresponding synthetic set for a training data size $K$.
For CIFAR10 and LSUN-CHURCH, we utilized the threshold $\delta_{d}$ as $0.03$ and $0.12$ respectively. We kept these threshold values similar to the memorized threshold values $\delta_{m}$ detailed in Table (2), since they have proven to be effective at separating the memorized samples from non-memorized ones in the previous experiment. However, since it is difficult to select an appropriate threshold value for the $L_{2}$-distance metric (due to its range being between 0 and $\infty$), we opted to use LPIPS to experiment with MNIST and FASHION-MNIST using the same threshold $\delta_{d}=0.1$ for both. To find this optimal threshold value, we selected three data sizes, where each belongs to one of the three phases: memorization, spurious, or generalization. We then observed which of the several threshold values, starting from $0.01$ to $0.15$ with the increment of $0.01$, reflects best of the performance given by LPIPS using vgg-16 [56] as its backbone. Specifically, we evaluated these thresholds using $K\in\{4000,11724,60000\}$ for MNIST, and $K\in\{4000,8965,60000\}$ for FASHION-MNIST. We simply performed Alg. (1) for those three data sizes on a small batch of $32$ samples for the three sample types, and manually checked whether the recovered samples (using the obtained critical time $t_{c}$ to denoise from) are similar to the original images and if this observation is consistent across those three data sizes.
C.3  Average Log-Volume Results
@@
Figure 16: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in CIFAR10 [37] and LSUN-CHURCH [75] using Eq. (34) with up to 512 samples per type. When fewer than 512 samples are available, all points are used. The average log-volume of training samples and the transition plot are included to confirm the shrinking basins of attraction in diffusion models. Variations in the critical time $t_{c}$ during the transition, which affect the radius $R$, lead to distinct log-volume trends across sample types, see the recorded critical times in Fig. (18). Shaded regions indicate standard deviation of the log-volume.
@@
Figure 17: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in FASHION-MNIST [73] and MNIST [17], computed using Eq. (34) with up to 512 samples per type. When fewer than 512 samples are available, all points are used. The average log-volume of training samples and the transition plot are included to confirm the shrinking basins of attraction in diffusion models. Variations in critical time $t_{c}$ and radius $R$ lead to distinct log-volume trends across sample types, see the recorded critical times in Fig. (19). Shaded regions indicate standard deviation of the log-volume.
C.4  Average Critical Time Results
@@
Figure 18: Average critical time, obtained using Alg. (1), with a maximum of 512 samples across the memorization-generalization transition for each sample type in CIFAR10 [37] and LSUN-CHURCH [75] The average critical times for the training data samples at each training data size $K$ are included to illustrate the overall decrease in the time and also the amount of perturbation in which the diffusion model can effectively recover the original sample within some error. Each sample type exhibits a distinctive critical time at each training data size $K$. The shaded region (or the error bars) shown here is the standard deviation computed from all of the critical times at each $K$ with respect to each type of samples. The transition plot of each dataset is shown here to visually guide where the memorization-generalization transition occurs.
@@
Figure 19: Average critical time, obtained using Alg. (1), with a maximum of 512 samples across the memorization-generalization transition for each sample type in FASHION-MNIST [73] and MNIST [17] The average critical times for the training data samples at each training data size $K$ are included to illustrate the overall decrease in the time and also the amount of perturbation in which the diffusion model can effectively recover the original sample within some error. Each sample type exhibits a distinctive critical time at each training data size $K$. The shaded region (or the error bars) shown here is the standard deviation computed from all of the critical times at each $K$ with respect to each type of samples. The transition plot of each dataset is shown here to visually guide where the memorization-generalization transition occurs.
Appendix D The Curvature of the Energy
D.1  Overview
Following [3, 66], we utilized their stable version of algorithm for computing the singular values of the Jacobian of the score function, detailed in Alg. (2). The motivations for the algorithm are first detailed in [63] and later expanded by [3, 66]. Here we recite the motivations laid out in [66].
Consider a generative diffusion model with $p_{0}({\mathbf{x}})$ defined on a $d$-dimensional manifold $\mathcal{M}$, where $\mathcal{M}=\mathcal{M}_{0}$ denotes the manifold at $t=0$. We can define a time-dependent set of points:
$\mathcal{M}_{t}=\{{\mathbf{x}}^{*}|\tilde{s}_{\mathcal{M}}({\mathbf{x}}^{*},t)=0,\,\,\text{with}\,\,J_{\mathcal{M}}({\mathbf{x}}^{*},t)\,\,\text{n.s.d.}\}$ (35)
which [66] name the stable latent set of the diffusion process. The negative semi-definiteness (n.s.d.) is a stability condition of the Jacobian matrix $J_{\mathcal{M}}({\mathbf{x}}^{*},t)$ of the support score $\tilde{s}_{\mathcal{M}}({\mathbf{x}}^{*},t)$, defined as the score function obtained from the uniform data distribution $\tilde{p}_{0}({\mathbf{x}})=\frac{1}{|\mathcal{M}|}\delta_{\mathcal{M}}({\mathbf{x}})$.
Using the relationship $s_{\theta}({\mathbf{x}},t)=-\frac{\epsilon_{\theta}({\mathbf{x}},t)}{\sigma_{t}}$, the diffusing particles typically explore shells of a radius that concentrates on $\sigma_{t}$. Then, for a small perturbation $\mathbf{p}$} around a state $\mathbf{x}$$^{*}$ on the latent manifold $\mathcal{M}_{t}$, the score function is well approximated by linearization:
$s({\mathbf{p}},t)\approx J({\mathbf{x}}^{*},t){\mathbf{p}}=\sum_{j}({\mathbf{v}}_{j}\cdot{\mathbf{p}})\lambda_{j}({\mathbf{x}}^{*},t){\mathbf{v}}_{j}$ (36)
where ${\mathbf{v}}_{j}$ and $\lambda_{j}$ are the respective $j$-th eigenvector and eigenvalue of the negative Jacobian $J({\mathbf{x}}^{*},t)$. Small perturbations, which aligned with the tangent space of $\mathcal{M}_{t}$, correspond to small eigenvalues; while orthogonal perturbations correspond to high eigenvalues, as the score tends to push the stochastic dynamics towards its fixed-points. Hence, based on these relationships, the spectrum of eigenvalues can provide detailed information regarding the local geometry of the stable latent set (35). Moreover, as already indicated by [63], one can estimate the dimensionality of the manifold from the location of a drop, or a sharp change, in the sorted spectrum of eigenvalues. The full range of the spectra results is shown in Figs. (20)-(23).
D.2  Experimental Details
Using Alg. (2), we computed the singular values of the Jacobian of the learnt score function $s_{\theta}$ given a small time $t_{0}$ for all datasets shown in Fig. (5A) of the main text. For each sample type, we utilized a maximum of 512 samples and computed the singular values at $t_{0}=2$ for all datasets and training data sizes. The singular values are sorted from greatest to least, and the top-5 values are ignored to avoid the case of very large singular values skewing the visualizations. Meanwhile, for the spectra results of the training data points shown in the bottom row of Fig. (5A), we utilized $\min(K,2048)$ number of training samples for all datasets and training data sizes. In this visualization, we utilized the rainbow color scheme and plotted the singular values for small training data sizes, starting from the purple color, to larger ones, ending at the red color.
Finally, regarding our experiment with Stable Diffusion [50], following [72, 30] we utilized its version 1.4 from the package Diffusers [69]. The candidate memorized and spurious prompts we utilized are originally found by [71, 70]. Additional examples are shown in Fig. (24). Meanwhile, for candidate generalized samples, their prompts are obtained from LAION2B-Aesthetic [54]. Instead of using Alg. (2), we opted for Alg. (3) and computed the singular values of the Jacobian of the score using $t_{0}=5$, due to the lack of computational resources. Please note that according to [66] Alg. (2) produces a much smaller number of extreme singular values than the later algorithm. For generation of the samples, we utilized DDIM [60] sampler with different noise vectors initialized from a fixed seed. For the visualizations, we skipped the top-50 singular values and set the limit for the y-axis as 1500.
Inputs
   Target sample ${\mathbf{x}}_{0}\in\mathbb{R}^{N}$
  2.5pt Time $t_{0}$
  2.5pt Score model $s_{\theta}$
  2.5pt Forward process $\mathcal{F}$
end
[…]
// Create an empty matrix
1exCompute matrix $\mathsf{M}$
 for $i=1,2,\dots,4N$ do
     5pt $\epsilon\sim\mathcal{N}(0,I)$
    // Sample noise
    ${\mathbf{x}}^{+}\leftarrow\mathcal{F}({\mathbf{x}}_{0},\epsilon,t_{0})$
    // Left perturbation
    ${\mathbf{x}}^{-}\leftarrow\mathcal{F}({\mathbf{x}}_{0},-\epsilon,t_{0})$
    // Right perturbation
    $\mathsf{M}^{(i)}\leftarrow\frac{s_{\theta}({\mathbf{x}}^{+},t_{0})-s_{\theta}({\mathbf{x}}^{-},t_{0})}{2}$
    // Add a new column to $\mathsf{M}$
   end for
end
Output
  5pt $\{s_{i}\}^{N}_{i=1}\leftarrow\text{SVD}(\mathsf{M})$
 // Compute singular values
end
Algorithm 2 Compute singular values of the Hessian of the energy using central difference for a pattern
Inputs
   Target sample ${\mathbf{x}}_{0}\in\mathbb{R}^{N}$
  2.5pt Time $t_{0}$
  2.5pt Score model $s_{\theta}$
  2.5pt Forward process $\mathcal{F}$
end
[…]
// Create an empty matrix
1exCompute matrix $\mathsf{M}$
 for $i=1,2,\dots,4N$ do
     5pt $\epsilon\sim\mathcal{N}(0,I)$
    // Sample noise
    ${\mathbf{x}}^{+}\leftarrow\mathcal{F}({\mathbf{x}}_{0},\epsilon,t_{0})$
    // Perturbation
    $\mathsf{M}^{(i)}\leftarrow s_{\theta}({\mathbf{x}}^{+},t_{0})$
    // Add a new column to $\mathsf{M}$
   end for
end
Output
  5pt $\{s_{i}\}^{N}_{i=1}\leftarrow\text{SVD}(\mathsf{M})$
 // Compute singular values
end
Algorithm 3 Compute singular values of the Hessian of the energy for a pattern
D.3  Curvature Results
Figure 20: Average singular value spectra of the score’s Jacobian for memorized, spurious, and generalized samples from diffusion models trained on the LSUN-CHURCH [75] dataset, shown across 38 increasing training data sizes $K$. The plots reveal a clear geometric hierarchy in the intermediate training regime, where memorized samples consistently exhibit the highest curvature (largest singular values), followed by spurious, and finally generalized samples. As the model enters the full generalization phase at large $K$, the average spectra for all three sample types become very distinguishable. Shaded region is the standard deviation of the singular values of different samples of each category.
Figure 21: Average singular value spectra of the score’s Jacobian for memorized, spurious, and generalized samples from diffusion models trained on the FASHION-MNIST [73] dataset, shown across 38 increasing training data sizes $K$. The plots reveal a clear geometric hierarchy in the intermediate training regime, where memorized samples consistently exhibit the highest curvature (largest singular values), followed by spurious, and finally generalized samples. As the model enters the full generalization phase at large $K$, the average spectra for all three sample types become very distinguishable. Shaded region is the standard deviation of the singular values of different samples of each category.
Figure 22: Average singular value spectra of the score’s Jacobian for memorized, spurious, and generalized samples from diffusion models trained on the MNIST [17] dataset, shown across 38 increasing training data sizes $K$. The plots reveal a clear geometric hierarchy in the intermediate training regime, where memorized samples consistently exhibit the highest curvature (largest singular values), followed by spurious, and finally generalized samples. As the model enters the full generalization phase at large $K$, the average spectra for all three sample types become very distinguishable. Shaded region is the standard deviation of the singular values of different samples of each category.
Figure 23: Average singular value spectra of the score’s Jacobian for memorized, spurious, and generalized samples from diffusion models trained on the CIFAR10 [37] dataset, shown across 38 increasing training data sizes $K$. In contrast to the clear separation as seen in other datasets, the spectra for all three sample types remain heavily overlapped throughout the entire transition from memorization to generalization. However, as the model enters the full generalization phase at large $K$, the average spectra for all three sample types become less indistinguishable. Shaded region is the standard deviation of the singular values of different samples of each category.
D.4  Additional Examples of Stable Diffusion
Figure 24: Additional examples of candidate memorized, spurious, and generalized samples from Stable Diffusion trained on LAION dataset [54], each corresponding to a distinct curvature signature. The candidate memorized sample has much larger singular values, while the candidate generalized sample has much smaller ones. The selected possible spurious samples have larger singular values than the candidate generalized sample, but much less in contrast to the candidate memorized sample. The common trait between candidate memorized and spurious samples is that both are stable attractors, as demonstrated by their repeated similar generations given different initial points (or noise vectors) and the conditioning on text-prompt. The y-axis is clipped at the value of 1500 to better contrast the shown examples’ spectra.
Appendix E Relative Energy Gap
E.1  Overview
In this section, we provide an additional experiment on the relative energy gap, which measures the gap between the energy depth of training data points and each of the three distinct sample types, using the method for approximating the relative potential from [49].
Consider the reverse SDE of the main text:
$\mathrm{d}{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})]\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}.$ (37)
We can re-express the above equation with the potential function $u(x,t)$:
$\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)=\mathbf{f}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ (38)